OpenStudy (anonymous):

∫ (2 - cosx + sinx) / (sin^2 x) dx

OpenStudy (underhill):

Just separate the integral into three parts, and solve each independently.

OpenStudy (anonymous):

Break it up into 3 integrals giving: \[2 \int\limits \frac{dx}{\sin^2(x)}-\int\limits \frac{\cos(x)}{\sin^2(x)}dx+\int\limits \frac{1}{\sin(x)}dx\]

OpenStudy (anonymous):

The answer is : \[-2\cot x + cosec x + \log (\sin x/2) - \log (\cos x/2) +c\]